Lang Industrial Systems Company (LISC) is trying to decide between two different conveyor belt systems. System A costs $232,000, has a four-year life, and requires $73,000 in pretax annual operating costs. System B costs $330,000, has a six-year life, and requires $67,000 in pretax annual operating costs. Both systems are to be depreciated straight-line to zero over their lives and will have zero salvage value. Whichever project is chosen, it will not be replaced when it wears out. The tax rate is 34 percent and the discount rate is 8 percent.

Calculate the NPV for both conveyor belt systems

NPV
  System A $   
  System B $   
Which conveyor belt system should the firm choose?
System B
System A
Answer #1
Calculation Of Post Tax Operating Cost
System A System A
a Pretax Operating Cost 73000 67000
b Add: Depreciation 58000 55000
c Subtotal 131000 122000
d Less: Saving In Tax 44540 41480
(c*34%)
e Less: Depreciation 58000 55000
f Operating Cost 28460 25520
(post tax)
Calculation Of NPV
Year Particular Discount Factor System A System B
@8% Cash Flow Discounted Cash Flow Cash Flow Discounted Cash Flow
a b c d e=d*c f g=f*c
0 Initial Cash Flow 1 232000 232000 330000 330000
1 Annual Operating Cost 0.926 28460 26352 25520 23630
2 Annual Operating Cost 0.857 28460 24400 25520 21879
3 Annual Operating Cost 0.794 28460 22592 25520 20259
4 Annual Operating Cost 0.735 28460 20919 25520 18758
5 Annual Operating Cost 0.681 25520 17368
6 Annual Operating Cost 0.630 25520 16082
NPV 326263 447976
Since NPV of System A is less than system B , the company should choose system A.
Note : Since it is not given in the question , answer is rounded off at nearest dollar without rounding off intermediate calculation.
Due to this there may be minor difference in answer.